The greatest term of the expansion (2x + 5y)13 where x = 10, y = 2 is
 
 
 
 
 
 
 
  None

 


 











 






























  The binomial expansion of n N contain a term independent of x :
 
   only if k is an integers
 
  only if k is a natural numbers.
 
  only if k is rational
 
  for any real k.

 


 











 






























  The coefficient of x50 in (1 + x)41 (1 - x + x2)40 is
 
  1
 
  2
 
  3
 
  0

 


 




























Given that the term of the expansion (x1/3 - x-1/2)15 which does not contain x is 5 m where m N , then m =
1100
1010
1001
none










0 out of 4 marks (content provider)

Correct Answer: 3
Solution:
Ans: 3





























Q. 8 In the binomial (21/3 + 3-1/3)n, if the ratio of the seventh term from the beginning of the expansion to the seventh term from its end is 1/6 , then n =
6
9
12
15










0 out of 4 marks (content provider)

Correct Answer: 2
Solution:
Ans: 2





























Q. 9 The coefficient of x49 in the expansion of (x - 1) ..... is equal to
- 2
+ ve coefficient of x
- ve coefficient of x
- 2










0 out of 4 marks (content provider)

Correct Answer: 1
Solution:
Ans: 1





























Q. 10 The last digit of (3P + 2) is :
where P = 34n and n N
1
2
4
5










0 out of 4 marks (content provider)

Correct Answer: 4
Solution:
Ans: 4





























Q. 11 The sum of the binomial coefficients of is equal to 256 . The constant term in the expansion is
1120
2110
1210
none










0 out of 4 marks (content provider)

Correct Answer: 1
Solution:
Ans: 1





























Q. 12 The coefficient of x4 in is :










0 out of 4 marks (content provider)

Correct Answer: 1
Solution:
Ans: 1





























Q. 13 If (1 + x + x2)25 = a0 + a1x + a2x2 + ..... + a50 . x50 then a0 + a2 + a4 + ..... + a50 is :
even
odd & of the form 3n
odd & of the form (3n - 1)
odd & of the form (3n + 1)










0 out of 4 marks (content provider)

Correct Answer: 1
Solution:
Ans: 1





























Q. 14 If the second term of the expansion is 14a5/2 then the value of is :
4
3
12
6










0 out of 4 marks (content provider)

Correct Answer: 1
Solution:
Ans: 1





























Q. 15 If (1 + x) (1 + x + x2) (1 + x + x2 + x3) ...... (1 + x + x2 + x3 + ...... + xn) a0 + a1x + a2x2 + a3x3 + ...... + amxm then has the value equal to
n!
(n + 1) !
(n - 1)!
none










0 out of 4 marks (content provider)

Correct Answer: 2
Solution:
Ans: 2





























Q. 16 The value of 4 {nC1 + 4 . nC2 + 42 . nC3 + ...... + 4n - 1} is :
0
5n + 1
5n
5n - 1










0 out of 4 marks (content provider)

Correct Answer: 4
Solution:
Ans: 4





























Q. 17 In the expansion of (1 + x)43 if the co-efficients of the (2r + 1)th and the (r + 2)th terms are equal, the value of r is :
12
13
14
15










0 out of 4 marks (content provider)

Correct Answer: 3
Solution:
Ans: 3





























Q. 18 The positive value of a so that the co-efficient of x5 is equal to that of x15 in the expansion of is
1
2










0 out of 4 marks (content provider)

Correct Answer: 1
Solution:
Ans: 1





























Q. 19 In the expansion of , the term which does not contain x is :
10C0
10C7
10C4
none










0 out of 4 marks (content provider)

Correct Answer: 3
Solution:
Ans: 3

















































Q. 20 Co-efficient of t in the expansion of, ( + p)m - 1 + ( + p)m - 2 ( + q) + ( + p)m - 3 ( + q)2 + ...... ( + q)m - 1
where - q and p q is :
 
 
 
 
 
 
 
 
 










0 out of 4 marks (content provider)

Correct Answer: 2
Solution:
Ans: 2

Grade:12th Pass

3 Answers

VISHWESH RAVI
13 years ago
answer soon. pls !!!!!!!!!!!!!! its urgent by VISHWESH RAVI
Badiuddin askIITians.ismu Expert
13 years ago

Dear Vishesh

1:

first calculate

  r= (n+1)/{1+(x/a)}

here n = 13 , x= 2x = 2*10=20

 a = 5y=5*2=10

so r = 14/3

integer value r=4

so Tr+1 will be the greatest term

i.e   T5

now T5 = 13C4 209 104

please post seperate question in seperate post

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin



 

VISHWESH RAVI
13 years ago

thanks a lot sir for the solution.

 

PLEASE PAY ATTENTION TO MY OTHER ASKED QUESTIONS (WITH TITLE PLEASE HELP).

I WILL BE VERY GRATEFUL TO YOU FOR THIS KIND STEP.

 

THANKING YOU,

VISHWESH RAVI